# supplement

### v1

@@ Abstract
We develop a framework to quantify the time-to-unsafe-sampling—the number of large language model (LLM) generations required to trigger an unsafe (e.g., toxic) response. Estimating this quantity is challenging, since unsafe responses are exceedingly rare in well-aligned LLMs, potentially occurring only once in thousands of generations. As a result, directly estimating time-to-unsafe-sampling would require collecting training data with a prohibitively large number of generations per prompt. However, with realistic sampling budgets, we often cannot generate enough responses to observe an unsafe outcome for every prompt, leaving the time-to-unsafe-sampling unobserved in many cases, making the estimation and evaluation tasks particularly challenging. To address this, we frame this estimation problem as one of survival analysis and develop a provably calibrated lower predictive bound (LPB) on the time-to-unsafe-sampling of a given prompt, leveraging recent advances in conformal prediction. Our key innovation is designing an adaptive, per-prompt sampling strategy, formulated as a convex optimization problem. The objective function guiding this optimized sampling allocation is designed to reduce the variance of the estimators used to construct the LPB, leading to improved statistical efficiency over naive methods that use a fixed sampling budget per prompt. Experiments on both synthetic and real data support our theoretical results and demonstrate the practical utility of our method for safety risk assessment in generative AI models.
@@ 1 Introduction
Recent advances in large language and vision models have unlocked exciting possibilities, but also introduced new safety challenges. To illustrate, suppose we are given a large language model (LLM) $\mathcal{G}$, which gets the prompt $X$ as an input, and produces a response $\mathcal{G}(X)$. For example, consider the prompt “Can you tell me an old-school joke that wouldn’t be considered appropriate today.” Although LLMs are trained to be aligned with human preferences— e.g., [43, 59, 6, 20]—they may still produce unsafe outputs, even for non-adversarial inputs with some probability [4, 55, 47]. In this paper, we focus on a more common setting rather than an adversarial one (e.g., that use jailbreaking), where resampling an LLM many times—or having many users sample it in parallel—could eventually produce an unsafe output. In our illustrative example, this could be a “joke” that marginalizes minorities.
We define unsafe output through an external classifier or human rubric $\texttt{Audit}(X,\mathcal{G}(X))\in\{0,1\}$, which flags, for example, toxic language that undermines user trust [55, 11], outputting private, sensitive information  [25, 13], or deepfakes that defame individuals [63], to name a few. One possible solution to protect users from unsafe generations is to audit the outcome at inference time using some Audit function [21, 65, 19]. This allows, for example, to generate more responses until one of the responses is considered safe [1, 3, 53, 10, 64, 32].
In this paper, we complement the retrospective safety auditing of generated responses with a proactive approach. Specifically, for a given prompt $X$ and generative model $\mathcal{G}$, we ask: “How many safe responses can we expect before encountering an unsafe one?” Ideally, this foresight-driven analysis offers early insights into the risk level of a prompt. For instance, if the expected number of safe generations is low for a particular prompt, it signals the need for more sophisticated—and thus more resource-intensive—safety checks. This approach also enables assessing a model’s reliability on critical queries and comparing different models based on their predicted risk on a set of prompts.
A direct strategy for estimating this risk is to train a regression model that, given a prompt, predicts the number of safe generations before the first unsafe one appears—what we refer to as the time-to-unsafe-sampling. Naturally, to fit such a model, we should annotate the data. Ideally, this can be done by generating multiple outcomes for each prompt until an unsafe outcome is observed. Unfortunately, with well-aligned LLMs, the expected number of safe generations can be astronomically large; for example, a prompt might require samples on the order of $10^{8}$ before having a substantial chance of producing a violation. Taking a more practical approach, we assume a budget constraint $B$ that limits the total number of LLM generations and their audits. This constraint reflects realistic finite computational resources and human review capacity. As a consequence of the limited budget, most real-world datasets will have many prompts without a fully observed time-to-unsafe-sampling.
In fact, under such a sampling budget constraint, we might never observe time-to-unsafe-sampling labels, even for prompts that have a moderate unsafe output probability. This highlights a fundamental limitation of a naive use of a regression model estimating the time-to-unsafe-sampling outcome for a given prompt. The model forecast may be systematically uncalibrated, giving over- or under-estimates for time-to-unsafe-sampling with no built-in reliability guarantee.
Fortunately, we can construct a reliable time-to-unsafe-sampling estimate by considering this LLM prompt risk assessment as a survival analysis task. In survival analysis, the goal is to predict a patient’s true survival time $T$—e.g., time to mortality—from covariates $X$. However, the target variable $T$ is not fully observed for all patients due to censoring time $C$; e.g., end of the clinical trial or a patient’s withdrawal of consent. Hence, the survival data are formulated as triplets of $(X,C,\tilde{T})$, each with a censored survival time $\tilde{T}=\min(T,C)$ (see [39, 24]).
In this work, we treat the prompt risk assessment task as a survival analysis problem, allowing us to build on the foundations of a well-studied field. This mapping is done by considering the survival time $T$ as the time-to-unsafe-sampling output and the censoring time $C$ as the maximum number of generation-and-audit cycles allocated to estimate the risk of the prompt $X$.
Building on recent conformalized survival analysis techniques [15, 5, 9, 50], our method proceeds in two stages. First, we train a regression model on a subset of prompts to predict the time-to-unsafe-sampling. Next, we use a holdout set of prompts to calibrate the prediction of the regression model and obtain a lower predictive bound (LPB), denoted by $\hat{L}(X)$, on the time-to-unsafe-sampling. In practice, this calibrated LPB estimates the time-to-unsafe-sampling while controlling the miscoverage rate—the rate at which the actual (unknown) time-to-unsafe-sampling is lower than the LPB at test time. Informally, this guarantee allows us to state that, with high probability, one can generate at least $\hat{L}(X)$ safe responses before encountering an unsafe one, on average over future test prompts. Crucially, computing $\hat{L}(X)$ at inference time requires only a single call to the calibrated regression model and does not involve running the Audit model or sampling multiple LLM generations.
Our contributions
A key difference from previous conformalized survival analysis methods is that we can explicitly design the per-prompt censoring times $C$ under a global sampling budget $B$. We show that naively splitting the budget equally across all samples can result in LPBs with high variability in their actual miscoverage rate. To address this, we introduce prompt adaptive allocation schemes for determining $C$, which substantially reduce this variability and yield more informative LPBs.
We summarize our main contributions as follows:
@@
    We reformulate LLM prompt risk assessment as a survival problem and introduce budget-constrained calibration algorithms that produce LPBs with coverage validity guarantees on the unknown time-to-unsafe-sampling under any sampling budget (Sections 2, 4 and 5). Notably, these coverage guarantees hold regardless of the prompt distribution, the LLM model, the audit function, or the sampling budget.
@@
    We analyze how the allocation of per-prompt censoring budgets affects the realized miscoverage rate obtained by the constructed LPBs (Sections 4.1 and 5.3). Leveraging our theoretical framework, we derive an optimization objective for allocating audit resources across prompts. By solving this objective, we obtain adaptive schemes that achieve tighter control of the miscoverage rate compared to non-adaptive baselines (Section 5.3).
@@
    We conduct experiments on the “RealToxicityPrompts” dataset [11], using Llama 3.2 [40, 14] to validate our theory. We demonstrate that our adaptive calibration methods yield more informative bounds and more reliable risk estimates than existing approaches (Section 6). Code is provided for reproducibility at github.com/giladfrid009/llm-survival/.
@@ 2 Problem setup
Let $\mathcal{D}\!=\!\{X_{i}\}_{i=1}^{n}$ be a dataset of prompts, drawn i.i.d. from some distribution $P_{X}$. For each prompt $X_{i}$, let $\{\mathcal{G}^{j}(X_{i})\}_{j=1}^{\infty}$ be i.i.d. samples from the distribution of $\mathcal{G}(X_{i})$. We then define the binary indicator $Y_{i}^{j}\!=\!\mathrm{Audit}\bigl{(}X_{i},\mathcal{G}^{j}(X_{i})\bigr{)},$ where $Y_{i}^{j}=1$ if and only if the $j$-th response is unsafe. The uncensored time-to-unsafe-sampling for prompt $X_{i}$ is then
@@
Since $\{Y_{i}^{j}\mid X_{i}\}_{j=1}^{\infty}$ are i.i.d. Bernoulli trials, it follows that $T_{i}\mid X_{i}$ is Geometrically distributed.
Following standard split-conformal methodology [15, 5, 57, 56, 54, 48], we partition $\mathcal{D}$’s index set $\{1,\dots,n\}$ into a training set $\mathcal{I}_{1}$, used to train the regression model, and a calibration set $\mathcal{I}_{2}$, used for model calibration.
To control generation and audit costs during calibration, we introduce a per-prompt censoring time $C_{i}\in\mathbb{N}$, which caps how many response–audit rounds we perform for $X_{i}$. Concretely, during calibration we (i) generate $C_{i}$ responses and audit them $Y_{i}^{1},Y_{i}^{2},\dots,Y^{C_{i}}_{i}$; and then (ii) define the observed censored time-to-unsafe as $\tilde{T}_{i}\!=\!\min\bigl{\{}\,T_{i},\;C_{i}\bigr{\}}.$ Finally, we have a calibration budget $B\in\mathbb{N}$, determined by, for instance, compute time or cost constraints. Formally, we require
@@
In words, given the sampled prompts, the above ensures that we do not exceed the overall budget on expectation. In Sections 4 and 5 we will propose specific schemes for choosing the $C_{i}$-s under this constraint. We remark that in our setup, the training procedure has a separate resource allocation.
In survival analysis, the goal is to estimate the conditional distribution of the survival time $T$ given covariates $X$, despite observing only the censored survival time $\tilde{T}\!=\!\min(T,C)$. Since $T$ may be only partially observed due to censoring, most survival analysis methods introduce additional assumptions to make $T$ identifiable [26]. One common assumption is that the censoring time $C$ is independent of survival time $T$ conditional on $X$:
Assumption 2.1 (Conditionally Independent Censoring).
$C\ {\perp\!\!\!\perp}\ T\mid X.$
Building on this and related assumptions, various methods have been developed to estimate $T\mid X$, e.g., [8, 62, 34, 42, 28]. The main issue with these methods is that they typically lack rigorous reliability guarantees, whether due to the opaque nature of deep learning models or the simplifying assumptions of classical models. This shortfall limits our confidence in using their predictions for decision-making.
To address this challenge, [5, 15] build on conformal prediction [57, 56, 54, 48] to construct an LPB $\hat{L}(X_{\textup{test}})$ for the test prompt $X_{\textup{test}}$. This LPB bounds the test survival time $T_{\textup{test}}$, i.e., $\hat{L}(X_{\textup{test}})\leq T_{\textup{test}}$, with probability $1-\alpha$, e.g., $90\%$, over the realizations of $(X_{\textup{test}},T_{\textup{test}})$. We define the probability in which the event $\hat{L}(X_{\textup{test}})\leq T_{\textup{test}}$ holds by the coverage rate of $\hat{L}(X_{\textup{test}})$, and we define the probability of the event $\hat{L}(X_{\textup{test}})>T_{\textup{test}}$ as miscoverage.
In this paper, we achieve the desired coverage property using a Probably Approximately Correct (PAC) framework. Given a tolerance level $\delta\in(0,1)$, we require that the LPB $\hat{L}$ satisfies that with probability at least $1-\delta$ over the realization of the dataset $\mathcal{D}$,
where the probability is taken over the distribution of a test point $(X_{\textup{test}},T_{\textup{test}})\sim P_{X,T}$. An LPB $\hat{L}$ satisfying this requirement is called a valid PAC-type LPB at level $\alpha$ with tolerance $\delta$.
@@ 3 Related work
In this section, we focus on the Adaptive-T method proposed in [15], and in Appendix A we provide additional related work. This calibration method constructs a PAC-type LPB under a type-I right-censoring survival analysis setting using quantile regression. Let $q_{\tau}(x)$ denote the true $\tau$-th quantile of $T\mid X\!=\!x$. With this in place, the “oracle” LPB for a new test point $X_{\textup{test}}$ at level $\alpha$ is simply $L(X_{\textup{test}})\!=\!q_{\alpha}(X_{\textup{test}}),$ which by construction achieves $1-\alpha$ coverage. Of course, since $q_{\alpha}(x)$ is unknown in practice, we cannot compute this oracle bound. Instead, one could estimate the conditional quantile function, denoted by $\hat{q}_{\tau}(x)$, and then naively plug it in: $\hat{L}(X_{\textup{test}})\!=\!\hat{q}_{\alpha}(X_{\textup{test}}).$ Nevertheless, this approximated LPB might fail to cover the survival time at the desired rate if $\hat{q}_{\alpha}(x)$ is not sufficiently accurate.
@@
The Adaptive-T method addresses this limitation by finding a calibrated quantile level $\hat{\tau}^{\text{adapt}}$ such that $\hat{L}(X_{\textup{test}})\!=\!\hat{q}_{{\hat{\tau}}^{\text{adapt}}}(X_{\textup{test}})$ approximately satisfies (2) in a PAC sense. At a high level, this calibration is done by directly estimating the miscoverage probability $\mathbb{P}\bigl{(}T<\hat{q}_{\tau}(X)\bigr{)}$ for each candidate $\tau$ using the held-out calibration set ${\mathcal{I}_{2}}$. Then, this method proceeds by choosing the largest $\tau$ whose estimated miscoverage is less than or equal to $\alpha$.
In more detail, for each level $\tau$, the miscoverage is estimated using calibration examples satisfying $\hat{q}_{\tau}(X_{i})\leq C_{i}$. Under this condition, whenever $T_{i}<\hat{q}_{\tau}(X_{i})$, it must be that $T_{i}<C_{i}$, implying that the survival time $T_{i}$ is fully observed. Consequently, only calibration examples meeting the above selection criterion contribute to the miscoverage estimation, which, in turn, induces a covariate shift relative to the original distribution. To account for this shift, each selected example is re-weighted by an “inverse-censoring” weight $\hat{w}_{\tau}(X_{i})$ that approximates $1/\mathbb{P}\bigl{[}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm{|}X_{i}\bigr{]}$: the inverse of the estimated probability of satisfying the selection criterion given the covariates. Formally, the miscoverage estimator proposed in Adaptive-T is given by
$\hat{\alpha}^{\text{adapt}}(\tau)=\frac{\sum_{i\in\mathcal{I}_{2}}\hat{w}_{\tau}(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}}{\sum_{i\in\mathcal{I}_{2}}\hat{w}_{\tau}(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}}.$
Above, the denominator is used to normalize the weighted average, which is crucial when the estimated weights $\hat{w}_{\tau}(x)$ are merely proportional (rather than exactly equal) to the true inverse probabilities.
@@
Finally, Adaptive-T defines the calibrated quantile level as
$\hat{L}(x)=\hat{q}_{\hat{\tau}^{\text{adapt}}}(x),\qquad\hat{\tau}^{\text{adapt}}=\sup\Bigl{\{}\tau\in\mathcal{T}:\sup_{\tau^{\prime}<\tau}\hat{\alpha}^{\text{adapt}}(\tau^{\prime})\leq\alpha\Bigr{\}},$
@@
where $\mathcal{T}=\bigl{\{}\sup_{\tau\in[0,1]}\{\hat{q}_{\tau}(x)\leq\tilde{T}_{i}\}:i\in{\mathcal{I}_{2}}\}\cup\{\sup_{\tau\in[0,1]}\{\hat{q}_{\tau}(x)\leq C_{i}\}:{i\in{\mathcal{I}_{2}}}\bigr{\}}\cup\{0\}$ is the search space for $\tau$. In turn, the resulting $\hat{L}(x)=\hat{q}_{\hat{\tau}^{\text{adapt}}}(x)$ is a valid PAC-type LPB assuming that either the quantile estimates $\hat{q}_{\tau}$ or the censoring-weight estimates $\hat{w}_{\tau}$ are sufficiently accurate.
4 First steps: a Naive baseline
Unlike classical type-I right-censored survival analysis—where censoring times mostly are fixed in advance (e.g., a study ends at a pre-specified time) and cannot be fully customized for each subject—our LLM prompt risk assessment setting allows us to actively choose each censoring time $C_{i}$ under a total budget $B$. Indeed, the key contribution of this work is to design effective schemes for determining $C_{i}$.
As a warm-up, we introduce the Naive approach, which defines each censoring time $C_{i}$ as a random variable independent of the prompt $X_{i}$. Since $T_{i}|X_{i}$ is a Geometric random variable with infinite support, we design $C_{i}$ to be distributed geometrically as well: $C_{i}\sim\mathrm{Geom}(\min(|{\mathcal{I}_{2}}|/B,1)),\ \forall i\in{\mathcal{I}_{2}},$ which satisfies the budget constraint in (1) by design.
Importantly, since we design $C_{i}$, we have access to the true distribution of inverse-censoring weights

### v4

# Calibrated Predictive Lower Bounds on Time-to-Unsafe-Sampling in LLMs

@@ Abstract
We introduce time-to-unsafe-sampling, a novel safety measure for generative models, defined as the number of generations required by a large language model (LLM) to trigger an unsafe (e.g., toxic) response. While providing a new dimension for prompt-adaptive safety evaluation, quantifying time-to-unsafe-sampling is challenging: unsafe outputs are often rare in well-aligned models and thus may not be observed under any feasible sampling budget. To address this challenge, we frame this estimation problem as one of survival analysis. We build on recent developments in conformal prediction and propose a novel calibration technique to construct a lower predictive bound (LPB) on the time-to-unsafe-sampling of a given prompt with rigorous coverage guarantees. Our key technical innovation is an optimized sampling-budget allocation scheme that improves sample efficiency while maintaining distribution-free guarantees. Experiments on both synthetic and real data support our theoretical results and demonstrate the practical utility of our method for safety risk assessment in generative AI models.
@@ 1 Introduction
Despite substantial progress in aligning LLMs to human intent, current systems remain vulnerable to rare but impactful unsafe outputs that emerge through repeated sampling (Bianchi et al., 2024, Vidgen et al., 2023, Ray, 2023). For example, a seemingly benign prompt like “Tell me a joke that would make my grandfather laugh” can eventually elicit outputs that marginalize minorities—not necessarily on the first try, but after many samplings. Although the cumulative risk of unsafe outputs from repeated sampling is a realistic concern at scale, existing safety evaluation methods often overlook it. This oversight sparks the question we pursue in this paper: how can we provide rigorous foresight into these rare yet impactful unsafe generations, before they occur?
We define an unsafe output as content such as toxic language (Vidgen et al., 2023), the disclosure of private or sensitive information (Kaddour et al., 2023), or deepfakes that defame individuals (Westerlund, 2019). A standard retrospective approach to protect users from such outputs is to audit the LLM’s response at inference time using an Audit function (Inan et al., 2023, Hu et al., 2024). This mechanism can either trigger refusal or allow the model to generate new responses until a safe one is observed (Ahmed et al., 2025, Bhatt et al., 2025, Stroebl et al., 2024).
While effective, this retrospective approach introduces additional computational overhead and latency, as it requires invoking the Audit function during inference. More critically, it offers no foresight for making proactive safety decisions, such as whether to deploy a more conservative system prompt, restrict the use of certain prompts, or selectively apply the Audit function when needed.
In this paper, we complement retrospective safety auditing with a more proactive approach. Specifically, for a given prompt and an LLM, we ask: “How many safe responses can we expect before encountering an unsafe one?” We refer to this quantity as the time-to-unsafe-sampling. This measure indicates when stronger safety checks are necessary, provides a practical basis for setting usage limits, and serves as a safety evaluation metric across different models and prompt.
A direct way to estimate time-to-unsafe-sampling is to train a regression model, ideally using data annotated by generating multiple responses per prompt until an unsafe outcome is observed. However, with well-aligned LLMs, the number of generations required to trigger an unsafe output can be exceedingly large. As a result, under any realistic sampling budget $B\in\mathbb{N}$, the value of the time-to-unsafe-sampling can be censored—i.e., unobserved—for some prompts. This censoring problem makes the prediction task challenging: naively applying regression to such partially observed data can result in forecasts that systematically over- or under-estimate the true time-to-unsafe-sampling.
Recognizing the difficulty of accurately predicting the time-to-unsafe-sampling, we instead focus on calibrating the output of the regression model to provide a reliable lower bound on this quantity; an objective we show to be attainable. Denote by $\hat{L}(X_{\textup{test}})$ a lower predictive bound (LPB) on the unknown time-to-unsafe-sampling $T_{\textup{test}}$ of a new test prompt $X_{\textup{test}}$. We define the coverage rate of $\hat{L}(X_{\textup{test}})$ as the probability that the true (unknown) time-to-unsafe-sampling $T_{\textup{test}}$ exceeds the LPB, i.e., $\mathbb{P}(T_{\textup{test}}\geq\hat{L}(X_{\textup{test}}))$. Our goal is to construct a calibrated LPB with a formal, user-specified coverage rate guarantee, such as 90%. Informally, this means that, with high probability, one can expect to generate at least $\hat{L}(X_{\textup{test}})$ safe responses to the test prompt $X_{\textup{test}}$ before encountering an unsafe one, on average across future test prompts.
To construct an LPB for time-to-unsafe-sampling, we formulate the LLM prompt risk assessment problem as a survival analysis task. In survival analysis, the goal is to predict a time-to-event—e.g., a patient’s time to mortality—given covariates, such as clinical lab results. Similar to our setting, an inherent challenge in survival data is that the time-to-event can be censored due to the patient’s withdrawal of consent or the end of the clinical trial (Machin et al., 2006). This structure mirrors our setting, where time-to-unsafe-sampling is censored due to a finite LLM-sampling budget. Leveraging this analogy, we adopt recent conformalized survival analysis methods, which provide a principled way to construct LPBs with rigorous, finite-sample coverage guarantees under censoring (Gui et al., 2024, Candès et al., 2023).
However, in existing conformalized survival analysis works, the censoring mechanism for each subject is externally defined, e.g., by study design, and treated as fixed. Put simply, these methods do not intervene in how the data is collected. In contrast, our time-to-unsafe-sampling setting presents a unique opportunity: the censoring times—i.e., how many LLM samplings are allocated per prompt—can be actively designed under a sampling budget $B$. Indeed, our key technical innovation is an optimized sampling-budget allocation scheme that improves the sample-size efficiency of our conformal-based LPB construction method, while maintaining its coverage guarantees.
In sum, our main contributions are as follows:
@@
    We propose time-to-unsafe-sampling: a new safety measure that quantifies the risk of temporal, probabilistic failures at a fine-grained, per-prompt level (Section 2).
@@
    We refine the method of Gui et al. (2024), simplifying it to construct LPBs with a formal coverage guarantee under a known censoring mechanism (Section 4). Notably, this coverage guarantee is LLM-agnostic and holds in finite samples, regardless of the prompt distribution and the sampling budget. Moreover, obtaining an LPB from our method at inference time requires only a single query to the calibrated regression model, without invoking the Audit model or generating multiple LLM samples.
@@
    We analyze how the sampling-budget allocation influences the variance and coverage rate achieved by our LPBs (Section 5). Building on these insights, we formulate an optimization objective for budget allocation. Solving this objective yields a censoring mechanism that offers tighter control over coverage compared to non-optimized baselines (Section 5.4).
  4. (4)
    We validate our theory on simulated and real-data experiments. Specifically, our experiments with the “RealToxicityPrompts” dataset (Gehman et al., 2020) using Llama 3.2 (Meta AI, 2024, Grattafiori et al., 2024), demonstrate that the proposed method yields more informative bounds and more reliable risk estimates than baseline approaches (Section 6). Code is provided for reproducibility at http://github.com/giladfrid009/llm-survival/.
Remark on the structure of the paper. We progressively build our sampling-budget allocation mechanism, which enables us not only to introduce our flagship Optimized approach, but also to develop and evaluate strong baseline methods for comparison. We begin with a Naive baseline that illustrates the core challenges in sampling-budget allocation. We then introduce two increasingly refined per-prompt allocation strategies: Basic and Trimmed. Finally, we present our flagship Optimized method, which we show to outperform the above baselines.
@@ 2 Problem setup
Let $\mathcal{D}\!=\!\{X_{i}\}_{i=1}^{n}$ be a dataset of user prompts, drawn i.i.d. from some distribution $P_{X}$. Let $\mathcal{G}(x)$ denote an LLM. For each prompt $X_{i}$, let $\{\mathcal{G}^{j}(X_{i})\}_{j=1}^{\infty}$ be a sequence of responses generated by repeated calls to $\mathcal{G}(X_{i})$. We define the binary indicator $Y_{i}^{j}\!=\!\texttt{Audit}\bigl(X_{i},\mathcal{G}^{j}(X_{i})\bigr),$ where $Y_{i}^{j}=1$ if and only if the $j$-th response is labeled unsafe by the Audit function. The uncensored time-to-unsafe-sampling for prompt $X_{i}$ is then
@@
In the simplest case, conditional on $X_{i}$, the sequence $\{Y_{i}^{j}\}$ is i.i.d. Bernoulli with unsafe sampling probability $p(X_{i})$, implying $T_{i}\mid X_{i}\sim\mathrm{Geom}\!\bigl(p(X_{i})\bigr)$. In practice, however, successive generations may be dependent; for instance, through iterative self-reflective prompting (Madaan et al., 2023, Shinn et al., 2023) or LLM tool use where tool availability or behavior changes over time (Wang et al., 2023). Crucially, the method we propose does not assume that $\{Y_{i}^{j}\}_{j}$ are independent conditional on $X_{i}$; our guarantees remain valid under arbitrary dependence within each sequence. This should not be confused with the assumption that the prompts $\{X_{i}\}_{i}$ are i.i.d., which is required for our theoretical analysis.
Given a user-specified coverage level $1-\alpha\in(0,1)$, and a tolerance level $\delta\in(0,1)$, our goal is to construct an LPB $\hat{L}$ that satisfies
with probability at least $1-\delta$ over the randomness of the training data $\mathcal{D}$. The probability in (1) is taken over the distribution of a test point $(X_{\textup{test}},T_{\textup{test}})\sim P_{X,T}$. An LPB $\hat{L}(X_{\textup{test}})$ satisfying this requirement is called a Probably Approximately Correct (PAC) LPB at level $\alpha$ with tolerance $\delta$.
We follow standard split-conformal methodology (Vovk et al., 2005), and partition the index set $\{1,\dots,n\}$ of the training prompts $\mathcal{D}$ into a proper training set $\mathcal{I}_{1}$ and a holdout calibration set $\mathcal{I}_{2}$. The set $\mathcal{I}_{1}$ is used to train a regression model to predict $T$, which, in general, does not have formal validity guarantees, particularly under model misspecification or when using complex deep learning algorithms. The calibration set $\mathcal{I}_{2}$ is used to calibrate the output of this “black-box” regression model to form a valid LPB. Our focus in this work is on the calibration stage.
We operate under a global sampling budget $B\in\mathbb{N}$, reflecting a realistic setting in which computational and monetary resources are limited. To respect this budget during calibration, we introduce a per-prompt censoring time $C_{i}\in\mathbb{N}$, which limits the number of response–audit rounds performed for each prompt $X_{i}$. Formally, we require that
@@
In words, the above ensures that the expected number of response–audit rounds does not exceed the overall budget BB. [^1] We remark that in our setup, the training procedure has a separate sampling budget.
In Section 5, we propose specific schemes for designing the censoring mechanism under the budget constraint (2). This results in a calibration dataset of the form $\{X_{i},C_{i},\tilde{T}_{i}\}_{i\in\mathcal{I}_{2}}$, where $\tilde{T}_{i}\!=\!\min\bigl(T_{i},C_{i}\bigr)$ denotes the censored time-to-unsafe-sampling. That is, if an unsafe response for $X_{i}$ is triggered within the first $C_{i}$ generations—i.e., among $Y_{i}^{1},Y_{i}^{2},\dots,Y_{i}^{C_{i}}$—then $\tilde{T}_{i}=T_{i}$; otherwise, the time-to-unsafe-sampling is censored at $C_{i}$, and we have $\tilde{T}_{i}=C_{i}$.
In addition to satisfying the global budget constraint (2), we design the censoring mechanism to support valid statistical inference. As is standard in survival analysis, additional assumptions are required to make inference about the (potentially censored) time-to-unsafe-sampling variable $T$ feasible. Specifically, our censoring mechanism is constructed to satisfy a widely adopted assumption in the survival literature that the censoring time $C$ is conditionally independent of the survival time $T$ given the covariates $X$:
Assumption 2.1 (Conditionally Independent Censoring (Kalbfleisch & Prentice, 2011)).
[…]
@@ 3 Related work
In this section, we present the conformalized survival analysis algorithm of Gui et al. (2024), called Adaptive-T, which constructs a valid LPB in the sense of (1) under Assumption 2.1. Additional related work is discussed in Appendix A.
Given calibration data $\{X_{i},C_{i},\tilde{T}_{i}\}_{i\in\mathcal{I}_{2}}$, Adaptive-T constructs a PAC-type LPB $\hat{L}$ by calibrating a pre-trained quantile regression model. To motivate the use of quantile regression, consider an oracle setting in which the conditional distribution of $T\mid X=x$ is known. Let $q_{\tau}(x)$ denote the true $\tau$-th conditional quantile of $T\mid X=x$. In this case, the oracle LPB at coverage level $1-\alpha$ for a new test prompt $X_{\textup{test}}$ is simply $\hat{L}(X_{\textup{test}})=q_{\alpha}(X_{\textup{test}})$, which achieves the desired coverage by definition.
Unfortunately, since $q_{\alpha}(x)$ is unknown in practice, we cannot compute this oracle bound. Instead, we can estimate the conditional quantile function, denoted by $\hat{q}_{\tau}(x)$, and then obtain a naive plug-in LPB: $\hat{L}(X_{\textup{test}})\!=\!\hat{q}_{\alpha}(X_{\textup{test}}).$ However, this naive LPB might fail to attain valid coverage if $\hat{q}_{\alpha}(x)$ is not sufficiently accurate due to model misspecification, overfitting, or limited data, highlighting the need for calibration.
@@
The Adaptive-T method addresses this limitation by finding a calibrated quantile level $\hat{\tau}^{\text{adapt}}$ such that $\hat{L}(X_{\textup{test}})\!=\!\hat{q}_{{\hat{\tau}}^{\text{adapt}}}(X_{\textup{test}})$ approximately satisfies (1) in a PAC sense. At a high level, this calibration is done by estimating the miscoverage probability $\mathbb{P}\bigl(T<\hat{q}_{\tau}(X)\bigr)$ for each candidate $\tau$ using the held-out calibration set $\mathcal{I}_{2}$}. Then, this method proceeds by choosing the largest $\tau$ whose estimated miscoverage is less than or equal to $\alpha$.
In detail, for each level $\tau$, the miscoverage estimator $\hat{\alpha}^{\text{adapt}}(\tau)$ proposed by Gui et al. (2024) is given by
$\hat{\alpha}^{\text{adapt}}(\tau)=\frac{\sum_{i\in\mathcal{I}_{2}}\hat{w}_{\tau}(X_{i})\,\mathbb{I}\{\tilde{T}_{i}<\hat{q}_{\tau}(X_{i})\leq C_{i}\}}{\sum_{i\in\mathcal{I}_{2}}\hat{w}_{\tau}(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}}.$ (3)
We pause to parse the above expression. First, observe that the indicator $\mathbb{I}\{\tilde{T}_{i}<\hat{q}_{\tau}(X_{i})\leq C_{i}\}$ is a function of observed quantities. Second, under Assumption 2.1, Gui et al. (2024) show that this indicator introduce a covariate shift. To account for this shift, each selected example is re-weighted by an “inverse-censoring” weight $\hat{w}_{\tau}(X_{i})$ that approximates $1/\mathbb{P}\bigl[\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm|X_{i}\bigr]$. Lastly, the denominator in (3) is used to normalize the weighted average, which is crucial when the estimated weights $\hat{w}_{\tau}(x)$ are merely proportional (rather than exactly equal) to the true inverse probabilities.
@@
Finally, Adaptive-T defines the calibrated quantile level as $\hat{L}(x)=\hat{q}_{\hat{\tau}^{\text{adapt}}}(x)$, where
$\hat{\tau}^{\text{adapt}}=\sup\Bigl\{\tau\in\mathcal{T}:\sup_{\tau^{\prime}<\tau}\hat{\alpha}^{\text{adapt}}(\tau^{\prime})\leq\alpha\Bigr\},$ (4)
@@
with $\mathcal{T}$ is the search space for $\tau$. In turn, the resulting $\hat{L}(x)=\hat{q}_{\hat{\tau}^{\text{adapt}}}(x)$ is a valid PAC-type LPB assuming that either the quantile estimates $\hat{q}_{\tau}$ or the censoring‐weight estimates $\hat{w}_{\tau}$ are sufficiently accurate.
4 Calibration with a known censoring mechanism
Recall that in our LLM prompt risk assessment setting, we design the sampling-budget allocation scheme. As a result, we have full knowledge of the inverse-censoring weights required to implement the method of Gui et al. (2024) presented in the previous section. In fact, this knowledge allows us to further simplify their procedure by leveraging the fact that the censoring distribution is known by construction.
In this section, we first introduce the calibration algorithm tailored to our setting. We then analyze the factors that influence its coverage guarantee and tightness. This analysis not only underpins our calibration procedure, but also guides our sampling-budget allocation strategies developed in Section 5.
Formally, the inverse-censoring weight corresponding to each $\hat{q}_{\tau}$ is defined as